# Survival Analysis with Adversarial Regularization

### v3

@@ Appendix A Appendix
Aids2 DRAFT 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
Noise 3.84e+05 1.61e+05 6.68e+04 2.78e+04 1.14e+04 4.42e+03 1.83e+03 8.54e+02 5.64e+02 5.20e+02 5.19e+02 5.21e+02
FGSM 6.77e+04 3.50e+04 1.78e+04 8.89e+03 4.35e+03 2.14e+03 1.10e+03 6.66e+02 5.35e+02 5.19e+02 5.19e+02 5.20e+02
PGD 5.49e+04 2.86e+04 1.49e+04 7.58e+03 3.79e+03 1.92e+03 1.02e+03 6.42e+02 5.31e+02 5.19e+02 5.19e+02 5.19e+02
SAWAR 5.29e+02 5.24e+02 5.22e+02 5.22e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.22e+02 5.22e+02 5.22e+02
Framingham DRAFT 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
Noise 1.27e+07 3.17e+06 8.03e+05 2.06e+05 5.38e+04 1.48e+04 4.92e+03 2.39e+03 1.72e+03 1.54e+03 1.51e+03 1.50e+03
FGSM 1.95e+06 5.96e+05 1.84e+05 5.70e+04 1.81e+04 6.33e+03 2.84e+03 1.85e+03 1.58e+03 1.51e+03 1.50e+03 1.49e+03
PGD 2.01e+07 4.36e+06 9.57e+05 2.13e+05 4.88e+04 1.21e+04 3.85e+03 1.97e+03 1.57e+03 1.50e+03 1.49e+03 1.49e+03
SAWAR 1.66e+03 1.62e+03 1.58e+03 1.56e+03 1.53e+03 1.52e+03 1.50e+03 1.50e+03 1.49e+03 1.49e+03 1.49e+03 1.49e+03
LeukSurv DRAFT 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
Noise 5.46e+23 1.79e+21 7.98e+18 2.52e+16 8.53e+13 1.88e+11 5.25e+08 2.35e+06 3.37e+04 2.60e+03 1.21e+03 7.21e+02
FGSM 8.34e+16 1.79e+15 3.10e+13 4.86e+11 8.13e+09 1.49e+08 2.72e+06 7.03e+04 3.45e+03 6.20e+02 4.31e+02 3.59e+02
PGD 1.28e+15 3.89e+13 1.03e+12 2.67e+10 7.96e+08 2.30e+07 7.01e+05 2.80e+04 1.98e+03 4.65e+02 3.58e+02 3.17e+02
SAWAR 2.07e+04 8.38e+03 3.50e+03 1.45e+03 6.62e+02 3.45e+02 2.91e+02 2.70e+02 2.60e+02 2.54e+02 2.52e+02 2.50e+02
TRACE DRAFT 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
Noise 9.11e+08 1.07e+08 1.34e+07 1.74e+06 2.43e+05 3.71e+04 6.63e+03 1.78e+03 8.35e+02 5.93e+02 5.51e+02 5.32e+02
FGSM 2.01e+11 9.10e+09 4.31e+08 2.24e+07 1.26e+06 8.26e+04 7.79e+03 1.61e+03 7.45e+02 5.74e+02 5.47e+02 5.35e+02
PGD 4.53e+08 4.81e+07 5.35e+06 6.12e+05 7.63e+04 1.07e+04 2.09e+03 8.37e+02 5.94e+02 5.40e+02 5.31e+02 5.26e+02
SAWAR 8.79e+02 7.60e+02 6.82e+02 6.29e+02 5.94e+02 5.69e+02 5.53e+02 5.41e+02 5.33e+02 5.28e+02 5.27e+02 5.27e+02
dataDIVAT1 DRAFT 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
Noise 5.69e+08 7.64e+07 1.05e+07 1.44e+06 2.07e+05 3.17e+04 5.75e+03 1.61e+03 9.08e+02 7.81e+02 7.67e+02 7.64e+02
FGSM 2.48e+08 3.36e+07 4.65e+06 6.56e+05 9.64e+04 1.54e+04 3.12e+03 1.14e+03 8.16e+02 7.65e+02 7.61e+02 7.60e+02
PGD 1.85e+11 8.02e+09 3.47e+08 1.75e+07 9.39e+05 6.14e+04 5.80e+03 1.27e+03 8.03e+02 7.57e+02 7.56e+02 7.57e+02
SAWAR 9.18e+02 8.55e+02 8.21e+02 8.00e+02 7.86e+02 7.75e+02 7.66e+02 7.59e+02 7.55e+02 7.52e+02 7.51e+02 7.51e+02
flchain DRAFT 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
Noise 2.18e+35 1.20e+31 7.58e+26 4.68e+22 3.48e+18 2.65e+14 2.21e+10 6.77e+06 4.45e+04 1.96e+03 1.22e+03 1.11e+03
FGSM 1.61e+20 3.47e+17 7.79e+14 1.86e+12 5.49e+09 4.02e+07 6.37e+05 1.86e+04 2.15e+03 1.15e+03 1.09e+03 1.08e+03
PGD 1.30e+16 7.91e+13 4.93e+11 3.47e+09 4.60e+07 1.33e+06 4.93e+04 4.99e+03 1.45e+03 1.10e+03 1.08e+03 1.07e+03
SAWAR 6.31e+03 2.22e+03 1.43e+03 1.19e+03 1.12e+03 1.09e+03 1.08e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03
prostate DRAFT 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
Noise 7.00e+20 5.05e+18 3.75e+16 2.85e+14 2.28e+12 2.03e+10 2.59e+08 9.04e+06 4.85e+05 6.66e+04 2.94e+04 1.37e+04
FGSM 3.16e+14 1.04e+13 3.40e+11 1.16e+10 4.07e+08 1.52e+07 6.47e+05 3.56e+04 3.75e+03 1.03e+03 7.20e+02 5.82e+02
PGD 8.84e+12 4.29e+11 2.18e+10 1.14e+09 5.94e+07 3.27e+06 1.92e+05 1.44e+04 1.91e+03 6.40e+02 5.00e+02 4.39e+02
SAWAR 9.65e+03 3.88e+03 1.66e+03 8.48e+02 5.43e+02 4.30e+02 3.83e+02 3.59e+02 3.45e+02 3.37e+02 3.35e+02 3.34e+02
retinopathy DRAFT 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
Noise 1.11e+04 5.61e+03 2.88e+03 1.51e+03 8.13e+02 4.68e+02 2.99e+02 2.19e+02 1.84e+02 1.70e+02 1.68e+02 1.68e+02
FGSM 7.43e+03 3.89e+03 2.07e+03 1.12e+03 6.33e+02 3.86e+02 2.63e+02 2.03e+02 1.78e+02 1.68e+02 1.67e+02 1.68e+02
PGD 7.30e+03 3.83e+03 2.04e+03 1.11e+03 6.25e+02 3.81e+02 2.60e+02 2.02e+02 1.77e+02 1.68e+02 1.67e+02 1.68e+02
SAWAR 4.02e+02 3.31e+02 2.79e+02 2.41e+02 2.14e+02 1.95e+02 1.82e+02 1.74e+02 1.69e+02 1.67e+02 1.67e+02 1.67e+02
stagec DRAFT 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
Noise 2.07e+06 4.93e+05 1.14e+05 2.73e+04 6.49e+03 1.61e+03 4.19e+02 1.25e+02 5.50e+01 4.10e+01 3.96e+01 3.95e+01
FGSM 1.80e+05 5.45e+04 1.67e+04 5.23e+03 1.64e+03 5.30e+02 1.77e+02 7.43e+01 4.54e+01 3.96e+01 3.91e+01 3.92e+01
PGD 1.54e+05 4.77e+04 1.47e+04 4.63e+03 1.47e+03 4.84e+02 1.69e+02 7.23e+01 4.48e+01 3.93e+01 3.89e+01 3.90e+01
SAWAR 1.03e+03 5.79e+02 3.30e+02 1.94e+02 1.21e+02 8.21e+01 6.10e+01 4.88e+01 4.22e+01 3.89e+01 3.83e+01 3.81e+01
zinc DRAFT 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
Noise 2.93e+07 4.72e+06 7.76e+05 1.32e+05 2.34e+04 4.45e+03 9.48e+02 2.56e+02 1.14e+02 8.17e+01 7.69e+01 7.53e+01
FGSM 2.50e+06 4.44e+05 9.49e+04 2.11e+04 4.89e+03 1.24e+03 3.66e+02 1.44e+02 9.01e+01 7.66e+01 7.53e+01 7.57e+01
PGD 2.55e+06 4.38e+05 8.27e+04 1.87e+04 4.43e+03 1.13e+03 3.40e+02 1.39e+02 8.88e+01 7.66e+01 7.55e+01 7.60e+01
SAWAR 1.31e+03 7.77e+02 4.72e+02 2.98e+02 1.98e+02 1.42e+02 1.10e+02 9.21e+01 8.28e+01 7.87e+01 7.80e+01 7.78e+01

### v4

@@ Appendix A Appendix
Dataset Algorithm
Aids2 DRAFT 7.88e+05 2.75e+05 9.79e+04 3.51e+04 1.31e+04 5.06e+03 2.06e+03 9.59e+02 6.16e+02 5.45e+02 5.41e+02 5.41e+02
Noise 3.15e+05 1.30e+05 5.45e+04 2.30e+04 9.53e+03 4.06e+03 1.79e+03 9.02e+02 6.06e+02 5.44e+02 5.41e+02 5.41e+02
FGSM 1.29e+05 5.99e+04 2.79e+04 1.29e+04 5.92e+03 2.72e+03 1.32e+03 7.49e+02 5.69e+02 5.42e+02 5.41e+02 5.41e+02
PGD 7.29e+04 3.62e+04 1.80e+04 8.83e+03 4.29e+03 2.10e+03 1.09e+03 6.82e+02 5.57e+02 5.42e+02 5.41e+02 5.41e+02
AAE-Cox 6.04e+12 2.43e+11 1.30e+10 6.94e+08 4.08e+07 2.64e+06 1.76e+05 1.27e+04 1.18e+03 5.41e+02 5.38e+02 5.39e+02
SAWAR 5.53e+02 5.48e+02 5.45e+02 5.43e+02 5.42e+02 5.42e+02 5.41e+02 5.41e+02 5.40e+02 5.40e+02 5.40e+02 5.40e+02
Framingham DRAFT 2.24e+07 4.61e+06 9.70e+05 2.13e+05 4.98e+04 1.33e+04 4.62e+03 2.35e+03 1.71e+03 1.52e+03 1.49e+03 1.48e+03
Noise 3.13e+07 6.71e+06 1.48e+06 3.35e+05 7.98e+04 2.05e+04 6.23e+03 2.67e+03 1.76e+03 1.53e+03 1.49e+03 1.48e+03
FGSM 8.34e+09 7.44e+08 6.79e+07 6.46e+06 6.50e+05 7.31e+04 1.05e+04 2.77e+03 1.67e+03 1.50e+03 1.49e+03 1.48e+03
PGD 1.53e+11 7.60e+09 3.88e+08 2.07e+07 1.20e+06 8.59e+04 9.43e+03 2.46e+03 1.61e+03 1.49e+03 1.48e+03 1.48e+03
AAE-Cox 1.38e+26 1.22e+23 1.20e+20 1.10e+17 1.21e+14 1.26e+11 1.63e+08 3.73e+05 3.68e+03 1.49e+03 1.47e+03 1.47e+03
SAWAR 1.67e+03 1.62e+03 1.58e+03 1.55e+03 1.53e+03 1.51e+03 1.50e+03 1.49e+03 1.48e+03 1.48e+03 1.48e+03 1.48e+03
LeukSurv DRAFT 2.10e+08 3.80e+07 7.18e+06 1.29e+06 2.41e+05 4.59e+04 9.21e+03 2.13e+03 6.86e+02 3.75e+02 3.28e+02 3.07e+02
Noise 1.40e+20 1.48e+18 1.45e+16 1.63e+14 1.29e+12 9.53e+09 8.92e+07 1.13e+06 2.90e+04 2.62e+03 1.25e+03 7.75e+02
FGSM 6.24e+12 4.01e+11 2.30e+10 1.27e+09 7.06e+07 4.34e+06 2.73e+05 2.04e+04 2.18e+03 5.11e+02 3.72e+02 3.25e+02
PGD 4.11e+11 3.30e+10 2.53e+09 1.99e+08 1.58e+07 1.30e+06 1.12e+05 1.10e+04 1.50e+03 4.34e+02 3.39e+02 3.07e+02
AAE-Cox 4.24e+28 2.71e+25 1.61e+34 7.46e+29 2.45e+25 7.92e+19 7.26e+14 3.55e+09 1.15e+05 4.18e+02 2.99e+02 2.84e+02
SAWAR 3.30e+03 2.00e+03 1.23e+03 7.95e+02 5.43e+02 3.97e+02 3.38e+02 3.06e+02 2.87e+02 2.75e+02 2.72e+02 2.70e+02
TRACE DRAFT 5.09e+05 1.75e+05 6.15e+04 2.23e+04 8.34e+03 3.31e+03 1.48e+03 8.34e+02 6.09e+02 5.38e+02 5.27e+02 5.24e+02
Noise 2.14e+09 2.12e+08 2.24e+07 2.58e+06 3.25e+05 4.62e+04 7.86e+03 1.89e+03 7.98e+02 5.76e+02 5.46e+02 5.35e+02
FGSM 4.68e+08 4.01e+07 3.85e+06 4.31e+05 5.82e+04 9.91e+03 2.30e+03 8.91e+02 6.03e+02 5.39e+02 5.31e+02 5.28e+02
PGD 4.42e+07 6.51e+06 1.04e+06 1.82e+05 3.49e+04 7.47e+03 2.01e+03 8.54e+02 5.99e+02 5.39e+02 5.31e+02 5.27e+02
AAE-Cox 1.06e+24 1.53e+21 2.08e+18 2.78e+15 2.40e+12 2.92e+09 5.29e+06 2.86e+04 9.75e+02 5.41e+02 5.31e+02 5.29e+02
SAWAR 1.49e+03 1.01e+03 7.74e+02 6.57e+02 5.98e+02 5.67e+02 5.48e+02 5.35e+02 5.28e+02 5.24e+02 5.24e+02 5.24e+02
dataDIVAT1 DRAFT 2.97e+04 1.56e+04 8.36e+03 4.56e+03 2.59e+03 1.59e+03 1.11e+03 8.83e+02 7.91e+02 7.56e+02 7.51e+02 7.50e+02
Noise 1.60e+10 1.12e+09 8.30e+07 6.62e+06 5.83e+05 5.99e+04 8.13e+03 1.87e+03 9.38e+02 7.78e+02 7.61e+02 7.59e+02
FGSM 3.25e+08 3.59e+07 4.13e+06 5.10e+05 7.07e+04 1.17e+04 2.66e+03 1.10e+03 8.11e+02 7.58e+02 7.53e+02 7.52e+02
PGD 1.22e+08 1.55e+07 2.06e+06 2.92e+05 4.56e+04 8.39e+03 2.15e+03 1.01e+03 7.94e+02 7.55e+02 7.52e+02 7.51e+02
AAE-Cox 9.96e+18 6.79e+16 3.12e+14 1.88e+12 1.37e+10 1.14e+08 1.05e+06 1.57e+04 1.04e+03 7.50e+02 7.45e+02 7.45e+02
SAWAR 7.89e+02 7.79e+02 7.70e+02 7.63e+02 7.58e+02 7.54e+02 7.51e+02 7.49e+02 7.47e+02 7.46e+02 7.46e+02 7.46e+02
flchain DRAFT 1.57e+22 5.59e+19 2.03e+17 7.97e+14 3.37e+12 1.52e+10 7.49e+07 5.25e+05 1.45e+04 1.95e+03 1.25e+03 1.10e+03
Noise 2.69e+32 2.90e+28 3.20e+24 5.94e+33 2.08e+27 7.50e+20 3.25e+14 2.31e+08 6.25e+04 2.40e+03 1.47e+03 1.24e+03
FGSM 5.19e+19 7.59e+16 1.21e+14 2.16e+11 7.32e+08 9.31e+06 2.85e+05 1.58e+04 2.16e+03 1.18e+03 1.12e+03 1.10e+03
PGD 3.10e+15 1.18e+13 6.11e+10 6.20e+08 1.63e+07 6.61e+05 3.81e+04 4.41e+03 1.43e+03 1.12e+03 1.09e+03 1.09e+03
AAE-Cox nan nan nan nan nan nan nan nan 4.02e+33 2.23e+10 1.11e+03 1.08e+03
SAWAR 5.84e+05 3.37e+04 4.04e+03 1.74e+03 1.31e+03 1.18e+03 1.13e+03 1.11e+03 1.10e+03 1.09e+03 1.09e+03 1.09e+03
prostate DRAFT 5.92e+05 2.18e+05 8.11e+04 3.04e+04 1.16e+04 4.49e+03 1.83e+03 8.39e+02 4.86e+02 3.71e+02 3.48e+02 3.37e+02
Noise 6.55e+23 1.63e+21 4.14e+18 1.02e+16 2.79e+13 8.22e+10 2.79e+08 1.36e+06 3.02e+04 3.29e+03 1.53e+03 9.01e+02
FGSM 5.24e+15 9.63e+13 1.84e+12 3.69e+10 7.62e+08 1.69e+07 4.35e+05 1.56e+04 1.19e+03 4.22e+02 3.76e+02 3.63e+02
PGD 1.20e+13 4.77e+11 1.96e+10 8.28e+08 3.66e+07 1.75e+06 9.45e+04 6.42e+03 8.39e+02 3.99e+02 3.67e+02 3.57e+02
AAE-Cox 1.53e+20 9.61e+17 6.16e+15 4.00e+13 2.65e+11 1.87e+09 1.33e+07 1.16e+05 1.81e+03 3.47e+02 3.33e+02 3.31e+02
SAWAR 6.21e+02 5.06e+02 4.28e+02 3.91e+02 3.67e+02 3.55e+02 3.47e+02 3.41e+02 3.37e+02 3.35e+02 3.34e+02 3.33e+02
retinopathy DRAFT 1.21e+04 6.19e+03 3.21e+03 1.68e+03 9.15e+02 5.24e+02 3.31e+02 2.39e+02 1.96e+02 1.76e+02 1.71e+02 1.69e+02
Noise 1.65e+04 8.10e+03 4.01e+03 2.03e+03 1.05e+03 5.83e+02 3.54e+02 2.47e+02 1.99e+02 1.77e+02 1.72e+02 1.69e+02
FGSM 7.62e+03 4.03e+03 2.15e+03 1.18e+03 6.72e+02 4.10e+02 2.80e+02 2.16e+02 1.86e+02 1.73e+02 1.70e+02 1.69e+02
PGD 7.32e+03 3.87e+03 2.08e+03 1.14e+03 6.55e+02 4.03e+02 2.75e+02 2.14e+02 1.85e+02 1.73e+02 1.70e+02 1.69e+02
AAE-Cox 1.32e+13 3.68e+11 1.03e+10 2.98e+08 8.67e+06 2.82e+05 1.17e+04 8.14e+02 1.98e+02 1.68e+02 1.67e+02 1.67e+02
SAWAR 6.10e+02 4.69e+02 3.69e+02 3.00e+02 2.51e+02 2.18e+02 1.96e+02 1.82e+02 1.74e+02 1.69e+02 1.68e+02 1.68e+02
stagec DRAFT 1.17e+04 5.05e+03 2.20e+03 9.79e+02 4.44e+02 2.10e+02 1.07e+02 6.47e+01 4.80e+01 4.24e+01 4.15e+01 4.13e+01
Noise 6.21e+04 2.17e+04 7.71e+03 2.72e+03 1.01e+03 3.86e+02 1.63e+02 8.11e+01 5.29e+01 4.44e+01 4.31e+01 4.27e+01
FGSM 1.02e+04 4.31e+03 1.80e+03 7.90e+02 3.54e+02 1.69e+02 9.18e+01 5.99e+01 4.77e+01 4.40e+01 4.35e+01 4.36e+01
PGD 8.55e+03 3.69e+03 1.60e+03 7.18e+02 3.32e+02 1.62e+02 8.89e+01 5.88e+01 4.73e+01 4.38e+01 4.34e+01 4.35e+01
AAE-Cox 1.32e+20 6.14e+17 2.58e+15 1.21e+13 5.37e+10 2.81e+08 1.53e+06 1.01e+04 1.53e+02 4.20e+01 4.10e+01 4.09e+01
SAWAR 2.52e+02 1.75e+02 1.26e+02 9.40e+01 7.32e+01 5.96e+01 5.14e+01 4.61e+01 4.28e+01 4.09e+01 4.04e+01 4.01e+01
zinc DRAFT 1.34e+06 3.43e+05 8.95e+04 2.36e+04 6.49e+03 1.84e+03 5.72e+02 2.12e+02 1.11e+02 8.33e+01 7.84e+01 7.64e+01
Noise 1.01e+07 2.02e+06 4.01e+05 8.52e+04 1.86e+04 4.26e+03 1.05e+03 3.06e+02 1.28e+02 8.63e+01 8.05e+01 7.86e+01
FGSM 3.49e+05 9.39e+04 2.69e+04 7.98e+03 2.42e+03 8.00e+02 3.01e+02 1.41e+02 9.24e+01 7.95e+01 7.82e+01 7.82e+01
PGD 4.28e+05 1.09e+05 2.89e+04 7.84e+03 2.37e+03 7.72e+02 2.87e+02 1.36e+02 9.09e+01 7.94e+01 7.83e+01 7.85e+01
AAE-Cox 1.15e+24 1.54e+21 4.97e+18 8.90e+15 1.88e+13 2.84e+10 5.64e+07 1.63e+05 7.39e+02 8.21e+01 7.79e+01 7.76e+01
SAWAR 5.06e+02 3.35e+02 2.31e+02 1.67e+02 1.28e+02 1.04e+02 9.02e+01 8.23e+01 7.82e+01 7.68e+01 7.68e+01 7.72e+01